# SVTRv2: CTC Beats Encoder-Decoder Models in Scene Text Recognition

### v1

@@ Abstract
Connectionist temporal classification (CTC)-based scene text recognition (STR) methods, e.g., SVTR, are widely employed in OCR applications, mainly due to their simple architecture, which only contains a visual model and a CTC-aligned linear classifier, and therefore fast inference. However, they generally have worse accuracy than encoder-decoder-based methods (EDTRs), particularly in challenging scenarios. In this paper, we propose SVTRv2, a CTC model that beats leading EDTRs in both accuracy and inference speed. SVTRv2 introduces novel upgrades to handle text irregularity and utilize linguistic context, which endows it with the capability to deal with challenging and diverse text instances. First, a multi-size resizing (MSR) strategy is proposed to adaptively resize the text and maintain its readability. Meanwhile, we introduce a feature rearrangement module (FRM) to ensure that visual features accommodate the alignment requirement of CTC well, thus alleviating the alignment puzzle. Second, we propose a semantic guidance module (SGM). It integrates linguistic context into the visual model, allowing it to leverage language information for improved accuracy. Moreover, SGM can be omitted at the inference stage and would not increase the inference cost. We evaluate SVTRv2 in both standard and recent challenging benchmarks, where SVTRv2 is fairly compared with 24 mainstream STR models across multiple scenarios, including different types of text irregularity, languages, and long text. The results indicate that SVTRv2 surpasses all the EDTRs across the scenarios in terms of accuracy and speed. Code is available at https://github.com/Topdu/OpenOCR.
@@ 1 Introduction
Typically, CTC-based methods [39, 11, 23, 28] employ a single visual model to extract image features and then apply a CTC-aligned linear classifier [16] to predict recognition results. This straightforward architecture provides advantages such as fast inference, which makes them especially popular in OCR applications. However, these models struggle to handle text irregularity, i.e., text distortions, varying layouts, etc. As a consequence, advanced attention-based decoders are introduced as alternatives to the CTC classifier, leading to a series of EDTRs [40, 38, 32, 29, 47, 55, 36, 59, 57, 15, 48, 4, 34, 46, 12, 58, 8, 62, 54, 49, 52, 9, 51, 17, 18, 63]. These attention-based decoders show appealing performance in integrating multi-modal cues, including visual [48, 58, 12, 52], linguistic [38, 55, 15, 36], and positional [57, 62, 8] ones, which are largely missed in current CTC models. This integration enables EDTRs to perform more effectively in complex scenarios. As depicted in the top of Fig. 1, compared to SVTR [11], a leading CTC model that is adopted by famous commercial OCR engines [28], EDTRs achieve superior results in English and Chinese benchmarks [25, 6], covering challenging scenarios such as curved, multi-oriented, occluded, and artistic text.
Nevertheless, EDTRs are commonly built upon complex architectures, thus sacrificing the inference speed (FPS), as shown in the top of Fig. 1. In addition, besides slower inference speed, EDTRs do not handle long text well. Even LISTER [8], an EDTR dedicated to long text recognition, performs worse than SVTR [11]. Since fast response and recognizing long text are both important for many applications, the OCR community has to face the dilemma that no model excels in accuracy, speed and versatility. When selecting either CTC-based models or EDTRs, users have to accept that the model is inferior in some aspects.
@@
In this paper, we aim to build more powerful CTC models by better handling text irregularity and integrating linguistic context. For the former, we address this challenge by first extracting discriminative features and then better aligning them. First, existing methods uniformly resize text images with various shapes to a fixed size before feeding into the visual model. We question the rationality of this resizing, which easily causes severe distortion of the text, making it unreadable to humans, as shown in the bottom-left of Fig. 1. To this end, we propose a multi-size resizing (MSR) strategy to adaptively resize text images according to their aspect ratios, thus minimizing text distortion and ensuring the discrimination of the extracted visual features. Second, irregular text may be rotated significantly, and the character arrangement does not align with the reading order of the text, causing the puzzle for CTC alignment, as shown in the bottom-centre example in Fig. 1. To solve this, we introduce a feature rearrangement module (FRM) which rearranges visual features with a horizontal rearrangement, and then identifying and prioritizing relevant vertical features. FRM maps 2D visual features into a sequence aligned with the text’ reading order, thus effectively alleviating the alignment puzzle. Consequently, CTC models integrating MSR and FRM can recognize irregular text effectively, without using rectification modules or attention-based decoders.
@@
As for the latter, the mistakenly recognized example shown in the bottom-right of Fig. 1 clearly highlights the necessity of integrating linguistic information. Since CTC models directly classify visual features, we have to endow the visual model with linguistic context modeling capability, which is less discussed previously. To this end, inspired by guided training of CTC (GTC) [23, 28] and string matching-based recognition [13], we propose a semantic guidance module (SGM), which devises a new scheme that solely leverages surrounding string context to recognize target characters during training. This approach effectively guides the visual model to learn to perceive the linguistic context without rely on decoders. During inference, SGM can be omitted and would not increase the inference cost.
@@
With these contributions, we develop SVTRv2, a novel CTC-based method whose recognition ability has been largely enhanced, while still maintaining a simple architecture and fast inference. To thoroughly validate SVTRv2, we conducted extensive ablation and comparative experiments on benchmarks including standard regular and irregular text [2], occluded scene text [48], the recent Union14M-L benchmarks [25], long text [13], and Chinese [6] text. The results demonstrate that SVTRv2 consistently outperforms all the compared EDTRs across the evaluated scenarios in terms of accuracy and speed, highlighting its effectiveness and broad applicability.
@@
In addition, recent advances [4, 25, 37] revealed the importance of large-scale real-world datasets in improving STR model performance. However, many STR methods primarily derived from synthetic data [19, 24], which fail to fully represent real-world complexities and lead to performance limitations, particularly on challenging scenarios. Notably, we observe that the existing real-word datasets [4, 25, 37] suffer from data leakage, and the results reported in [25] should be updated. As a result, we introduce U14M-Filter, a rigorously filtered version of the real-world training dataset Union14M-L [25]. We systematically reproduced and retrained 24 mainstream STR methods from scratch based on U14M-Filter. These methods are also thoroughly evaluated across various STR benchmarks. Their accuracy, model size, and inference time constitute a comprehensive and reliable new benchmark for future reference.
@@ 2 Related Work
Irregular text recognition [1, 35, 25] has posed a significant challenge in STR due to the diverse variation of text instances, where CTC-based methods [39, 11, 23, 28] are often less effective. To address this, some methods [40, 36, 59, 62, 54, 11, 61] incorporate rectification modules [40, 32, 61] that aim to transform irregular text into more regular format. Alternatively, more methods utilize attention-based decoders [47, 29, 38], which employ the attention mechanism to dynamically localize characters regardless of text layout, and thus less affected. However, these methods generally have tailored training hyper-parameters. For instance, the rectification modules [40, 32, 61] typically specify a fixed output image size (e.g. 32128), which is not always a suitable choice. While attention-based decoders [47, 29, 38] generally set the maximum recognition length to 25 characters, thus longer text cannot be correctly recognized, as shown in Fig. 5.
@@
Linguistic Context Modeling. There are several ways of modeling linguistic context. One major branch is auto-regressive (AR)-based STR methods [40, 47, 38, 29, 57, 25, 51, 62, 52, 54, 63, 14], which utilize previously decoded characters to model contextual cues. However, their inference speed is slow due to the character-by-character decoding nature. Some other methods [55, 15, 34, 4] integrate external language models to model the linguistic context and correct the recognition results. While effective, the linguistic context is purely text-based, making it challenging to adapt them to CTC models. There also are some studies [36, 48, 58] to modeling linguistic context with visual information only by using masked image modeling-based pretraining [21, 3]. However, they still depend on attention-based decoders to unleash the linguistic information, not integrating linguistic cues into the visual model, thus limiting their effectiveness in enhancing CTC models.
3 Methods
@@
Fig. 2 illustrates the overview of SVTRv2. A text image is first resized by MSR to the closest aspect ratio, forming the input X . X then experiences three consecutive feature extraction stages, yielding visual features F . During training, F is fed into both SGM and FRM. SGM guides SVTRv2 to model linguistic context, while FRM rearranges F into the character feature sequence , which is synchronized with the text reading order and aligns with the label sequence well. During inference, the SGM is discarded for efficiency.
3.1 Multi-Scale Resizing
@@
Previous works typically resize irregular text images to a fixed size, such as , which, however, may cause undesired text distortion and severely affect the quality of extracted visual features. To address this issue, we propose a simple yet effective multi-size resizing (MSR) strategy that resizes text shapes based on the aspect ratio (R=). Specifically, we define four specific sizes: [64, 64], [48, 96], [40, 112], and [32, 32], respectively corresponding to aspect ratio: R1.5 (R1), 1.5 R2.5 (R2), 2.5R3.5 (R3), and R3.5 (R4). Therefore, MSR allows text instances adaptively resized under the principles of roughly maintaining their aspect ratios, such that significant text distortion caused by resizing is almost eliminated. As a result, the quality of extracted visual features is guaranteed.
@@ 3.2 Visual Feature Extraction
Motivated by SVTR [11], the network architecture of SVTRv2 comprises three stages, with containing mixing blocks, as illustrated in Fig. 2. To extract discriminative visual features, we devise two types of mixing blocks: local and global. Local mixing is implemented through two consecutive grouped convolutions, which are expected to capture local character features, such as edges, textures, and strokes. Meanwhile, global mixing is realized by the multi-head self-attention (MHSA) mechanism [42]. This mechanism performs global contextual modeling on features, thereby enhancing the model’s comprehension of inter-character relationships and the overall text image. Both the number of groups in the grouped convolution and the number of heads in MHSA are set to . Similar to SVTR, by adjusting hyper-parameters and , we can derive three variants of SVTRv2 with different capacities, i.e., Tiny, Small, and Base, which are detailed in Sec. 7 of Supplementary.
3.3 Feature Rearranging Module
We propose a feature rearrangement module (FRM) to tackle the CTC alignment puzzle arising from rotated text. It rearranges the 2D features F into a feature sequence synchronized with the reading order of the text image. We regard this process as mapping the relevant features from to , where and . This this rearrangement can be formalized by a matrix , whereby can be derived from .
Grounding in the fact that the degree of text curve and rotation can be decomposed into offset components in both horizontal and vertical directions, we propose to learn M by using two distinct steps: horizontal rearrangement and vertical rearrangement. As described in Eq. 1, the horizontal one operates on , each row in the feature map, to learn a horizontal rearrangement matrix , where element represents the probability that the horizontal rearranged feature F corresponds to the original feature Fi,m. Based on the learned , the features in each row are rearranged on the horizontal direction. Subsequently, through a residual and MLP processing, we obtain F.
where are learnable weights, is the Softmax function, and F.
In the following vertical rearrangement, we introduce a selecting token, denoted as , which simultaneously attends to each column of features F within Fh to learn a vertical rearrangement matrix , as detailed in Eq. 2. The element represents the probability that the vertical rearranged feature F corresponds to F. Moreover, all column features share a single selecting token, rather than assigning a unique selecting token to each column feature at different locations. This scheme allows the model to generalize to longer text sequences, even when the number of column features exceeds the number seen during training, thereby facilitating the effective recognition of long text.
@@ 3.3 Feature Rearranging Module
We denote F as the rearranged feature sequence . Then, the predicted character sequence is obtained after passes through the classifier , and further aligned with the label sequence using the CTC rule, where is the learnable weight of the classifier. Furthermore, to intuitively clarify the role of FRM, we rewrite the mapping relation between and the original feature F by the following procedure:
| M | ||||
As seen in Eq. 4, the visual features F, which fully describe the text to be recognized but may be irregularly organized, are mapped to a more canonical . Different from existing text rectification models [40, 32], which apply geometric transformations at image space and often fail to correct severely irregular text, our FRM is carried out at feature space. It learns a condensed feature sequence by mapping and rearranging the relevant cues from F to proper positions in . By employing a structured way that the features are first rearranged horizontally and then vertically, FRM can accommodate text instances with significant irregularities such as large orientation variations, and obtaining features better aligned with the CTC classification.
3.4 Semantic Guidance Module
@@
CTC models directly classify visual features to obtain recognition results. This scheme inherently requires that the linguistic context must be incorporated into visual features, so that the CTC could benefit from. In light of this, we propose a semantic guidance module (SGM) as follows.
@@
For each character in a text image with character labels Y = {, , , }, we define its contextual information as the surrounding left string = and right string = , where denotes the context window length. The SGM’s role is to guide the visual model to integrate context from both and into visual features.
@@
We describe the process using the left string , with the same process applied symmetrically to the right string . Firstly, the characters in are mapped to string embeddings . Subsequently, these embeddings are encoded to create a hidden representation , representing the context of the left-side string . The attention map is computed by applying a dot product between the hidden representation and the visual features F, transformed by learned weight matrices and . The detailed formulation is as follows:
@@
where represents a pre-defined token that encodes the left-side string. The attention map is used to weight the visual features F, producing the feature corresponding to character . After processing through the classifier , the predicted class probabilities for is obtained to calculate the cross-entropy loss with the label , where is learnable weights of the classifier.
@@
The weight of the attention map records the relevance of to the visual features F, and moreover, represents the context of string . So only when the visual model incorporates the context from into the visual features of the target character , the attention map can maximize the relevance between and visual features of the character , thereby accurately highlighting the corresponding position of character , as shown in Fig. 3. A similar process applies to the right-side string , where the corresponding attention map and visual feature contribute to the prediction . By leveraging the above scheme during training, SGM effectively guides the visual model in integrating the linguistic context into visual features. Consequently, even when SGM is not used during inference, the linguistic context can still be maintained alongside the visual features, and enhancing the accuracy of CTC models. In contrast, previous methods, such as VisionLAN [48] and LPV [58], despite modeling linguistic context using visual features, still rely on attention-based decoders to unleash linguistic information, a process that is incompatible with CTC models.

### v2

@@ Abstract
Connectionist temporal classification (CTC)-based scene text recognition (STR) methods, e.g., SVTR, are widely employed in OCR applications, mainly due to their simple architecture, which only contains a visual model and a CTC-aligned linear classifier, and therefore fast inference. However, they generally exhibit worse accuracy than encoder-decoder-based methods (EDTRs) due to struggling with text irregularity and linguistic missing. To address these challenges, we propose SVTRv2, a CTC model endowed with the ability to handle text irregularities and model linguistic context. First, a multi-size resizing strategy is proposed to resize text instances to appropriate predefined sizes, effectively avoiding severe text distortion. Meanwhile, we introduce a feature rearrangement module to ensure that visual features accommodate the requirement of CTC, thus alleviating the alignment puzzle. Second, we propose a semantic guidance module. It integrates linguistic context into the visual features, allowing CTC model to leverage language information for accuracy improvement. This module can be omitted at the inference stage and would not increase the time cost. We extensively evaluate SVTRv2 in both standard and recent challenging benchmarks, where SVTRv2 is fairly compared to popular STR models across multiple scenarios, including different types of text irregularity, languages, long text, and whether employing pretraining. SVTRv2 surpasses most EDTRs across the scenarios in terms of accuracy and inference speed. Code: https://github.com/Topdu/OpenOCR.
@@ 1 Introduction
Typically, CTC-based methods [39, 11, 23, 28] employ a single visual model to extract image features and then apply a CTC-aligned linear classifier [17] to predict recognition results. This straightforward architecture provides advantages such as fast inference, which makes them especially popular in OCR applications. However, these models struggle to handle text irregularity, i.e., text distortions, varying layouts, etc. As a consequence, attention-based decoders are introduced as alternatives, leading to a series of EDTRs [40, 38, 32, 29, 46, 55, 36, 59, 57, 15, 47, 48, 4, 45, 13, 58, 8, 65, 16, 54, 49, 52, 9, 51, 18, 19, 67, 63, 66, 62, 61]. These methods exhibit superior performance in complex scenarios by leveraging multi-modal cues, including visual [47, 58, 13, 52], linguistic [38, 55, 15, 36], and positional [57, 65, 8] ones, which are largely missed in current CTC models. As depicted in the top of Fig. 1, compared to SVTR [11], a leading CTC model adopted by famous commercial OCR engines [28], EDTRs achieve superior results in scenarios [47, 25, 6] such as curved, multi-oriented, artistic, occluded, and Chinese text.
@@
In this paper, our aim is to build more powerful CTC models by better handling text irregularity and integrating linguistic context. For the former, we address this challenge by first extracting discriminative features and then better aligning them. First, existing methods uniformly resize text images with various shapes to a fixed size before feeding into the visual model. We question the rationality of this resizing, which easily causes unnecessary text distortion, making the text difficult to read, as shown in the bottom-left of Fig. 1. To this end, a multi-size resizing (MSR) strategy is proposed to resize the text instance to a proper predefined size based on its aspect ratio, thus minimizing text distortion and ensuring the discrimination of the extracted visual features. Second, irregular text may be rotated significantly, and the character arrangement does not align with the reading order of the text, causing the puzzle of CTC alignment, as shown in the bottom-center example in Fig. 1. To solve this, we introduce a feature rearrangement module (FRM). It rearranges visual features with first a horizontal rearrangement and then a vertical rearrangement to identify and prioritize relevant features. FRM maps 2D visual features into a sequence aligned with the text’s reading order, thus effectively alleviating the alignment puzzle. Consequently, CTC models integrating MSR and FRM can recognize irregular text well, without using rectification modules or attention-based decoders.
@@
As for the latter, the mistakenly recognized example shown in the bottom-right of Fig. 1 clearly highlights the necessity of integrating linguistic information. Since CTC models directly classify visual features, we have to endow the visual model with linguistic context modeling capability, which is less discussed previously. Inspired by guided training of CTC (GTC) [23, 28] and string matching-based recognition [12], we propose a semantic guidance module (SGM), a new scheme that solely leverages surrounding string context to model the target character. This approach effectively guides the visual model in capturing linguistic context. During inference, SGM can be omitted and would not increase the time cost.
@@
With these contributions, we develop SVTRv2, a novel CTC-based method whose recognition ability has been largely enhanced, while still maintaining a simple inference architecture and fast speed. To thoroughly validate SVTRv2, we conducted extensive and comparative experiments on benchmarks including standard regular and irregular text [2], Union14M-Benchmark [25], occluded scene text [47], long text [12], and Chinese text [6]. The results demonstrate that SVTRv2 consistently outperforms all the compared EDTRs across the evaluated scenarios in terms of accuracy and speed. Moreover, a simple pretraining on SVTRv2 yields highly competitive accuracy compared to the pretraining-based EDTRs advances [63, 66, 62, 61], highlighting its effectiveness and broad applicability.
@@
In addition, recent advances [25, 37] indicated the importance of large-scale real-world datasets in improving STR performance. However, many STR models primarily derived from synthetic data [20, 24], which fail to fully represent real-world complexities and lead to performance limitations, particularly on challenging scenarios. Meanwhile, we observe that existing large-scale real-word training datasets [4, 25, 37] overlap with Union14M-Benchmark, causing a small overlapping between training and test data, thus the results reported in [25] should be updated. As a result, we introduce U14M-Filter, a rigorously filtered version of the real-world training dataset Union14M-L [25]. Then, we systematically reproduced and retrained 24 mainstream STR methods from scratch based on U14M-Filter. These methods are thoroughly evaluated on Union14M-Benchmark. Their accuracy, model size, and inference time constitute a comprehensive and reliable new benchmark for future reference.
@@ 2 Related Work
Irregular text recognition [1, 35, 25] has posed a significant challenge in STR due to the diverse variation of text instances, where CTC-based methods [39, 11, 23, 28] are often less effective. To address this, some methods [40, 36, 59, 65, 54, 11, 64] incorporate rectification modules [40, 32, 64] that aim to transform irregular text into more regular format. While more methods utilize attention-based decoders [46, 29, 38, 57, 13, 12], which employ the attention mechanism to dynamically localize characters regardless of text layout, and thus are less affected. However, these methods generally have tailored training hyper-parameters. For example, rectification modules [40, 32, 64] typically specify a fixed output image size (e.g. 32128), which is not always a suitable choice. While attention-based decoders [46, 29, 38, 57, 13, 12] generally set the maximum recognition length to 25 characters, therefore, longer text cannot be recognized, as shown in Fig. 5.
@@
Linguistic context modeling. There are several ways of modeling linguistic context. One major branch is auto-regressive (AR)-based STR methods [40, 46, 38, 29, 25, 51, 65, 52, 54, 67, 14], which utilize previously decoded characters to model contextual cues. However, their inference speed is slow due to the character-by-character decoding nature. Some other methods [55, 15, 34, 4] integrate external language models to model linguistic context and correct the recognition results. While effective, the linguistic context is purely text-based, making it challenging to adapt them to the visual model of CTC models. There are also some studies [36, 47, 58] to model linguistic context with visual information only using pretraining based on masked image modeling [22, 3]. However, they still depend on attention-based decoders to utilize linguistic information, not integrating linguistic cues into the visual model, thus limiting their effectiveness in enhancing CTC models.
3 Method
@@
Fig. 2 illustrates the overview of SVTRv2. A text image is first resized by MSR to the closest aspect ratio, forming the input , which then experiences three consecutive feature extraction stages, yielding visual features . During training, is fed into both SGM and FRM. SGM guides SVTRv2 to model linguistic context, while FRM rearranges into the character feature sequence , which is synchronized with the text reading order and aligns with the label sequence. During inference, the SGM is discarded for efficiency.
3.1 Multi-Size Resizing (MSR)
@@
Previous works typically resize irregular text images to a fixed size, such as , which may cause undesired text distortion and severely affect the quality of extracted visual features. To address this issue, we propose a simple yet effective multi-size resizing (MSR) strategy that resizes text shapes based on the aspect ratio (). Specifically, we define four specific sizes: [64, 64], [48, 96], [40, 112], and [32, 32], respectively corresponding to aspect ratio: 1.5 (), 1.5 2.5 (), 2.5 3.5 (), and 3.5 (). Note that the first three buckets are fixed thus text instances in the same one can be trained in batch, while the fourth one can handle long text without introducing significant distortion. Therefore, MSR allows text instances adaptively resized under the principles of roughly maintaining their aspect ratios, and significant text distortion caused by resizing is almost eliminated.
@@ 3.2 Visual Feature Extraction
Motivated by SVTR [11], the visual model of SVTRv2 comprises three stages, with containing mixing blocks, as illustrated in Fig. 2. To extract discriminative visual features, we devise two types of mixing blocks: local and global. Unlike SVTR, for being able to handle multiple sizes, we do not use absolute positional encoding. In contrast, to model positional information, we implement local mixing as two consecutive grouped convolutions alternative to window attention [11], and effectively capturing local character features, such as edges, textures, and strokes. Meanwhile, global mixing is realized by the multi-head self-attention (MHSA) mechanism [41]. This mechanism performs global contextual modeling on features, thereby enhancing the model’s comprehension of inter-character relationships and the overall text image. Both the number of groups in the grouped convolution and the number of heads in MHSA are set to . Similar to SVTR [11], by adjusting hyper-parameters and , we derive three variants of SVTRv2 with different capacities, i.e., Tiny (T), Small (S), and Base (B), which are detailed in Suppl. Sec. 7.
3.3 Feature Rearranging Module (FRM)
To address the alignment puzzle caused by text irregularities like rotated text, we propose a feature rearrangement module (FRM). This module rearranges the visual features into a sequence that conforms to the CTC alignment requirement. We model the rearrangement process as a soft mapping using a probability matrix , where each element represents the probability of mapping the original feature to the rearranged feature , where and . Consequently, the softly rearrangement process is formalized as:
FRM is responsible for learning the matrix . To ensure that the module is sensitive to text orientation, we decompose the learning process of into two sequential steps: horizontal rearranging and vertical rearranging. As illustrated in Fig. 2, the horizontal one processes each row of the visual feature , denoted as , to learn a horizontal rearrangement matrix , which rearranges visual features along the horizontal direction. We implement this process using a multi-head self-attention mechanism as follows:
where are learnable weights, and denote the softmax function and matrix transpose operation, respectively. LN ans MLP means Layer Normalization and Multi-Layer Perceptron with an expansion rate of 4, respectively. represents the horizontally rearranged visual features.
Similarly, the vertical rearrangement processes visual features column-wise. Unlike the horizontal step, we introduce a selecting token , which interacts with all column features via cross-attention to learn a vertical rearrangement matrix . The elements of represent the probability of mapping column features to the rearranged features, yielding as follows:
@@ 3.3 Feature Rearranging Module (FRM)
We denote as the final rearranged feature sequence . The predicted character sequence is then obtained by passing through the classifier where is the learnable weight matrix. The predicted sequence is then aligned with the ground truth sequence using the CTC alignment rule.
3.4 Semantic Guidance Module (SGM)
@@
CTC models classify visual features directly to obtain recognition results. This scheme inherently requires that the linguistic context must be incorporated into visual features, only that the CTC could benefit from it. In light of this, we propose a semantic guidance module (SGM) as follows.
@@
For a text image with character labels = {, , , }, where is the -th character, we define its contextual information as the surrounding left string = and right string = , where denotes the size of the context window. SGM’s role is to guide the visual model to integrate context from both and into visual features.
@@
We describe the process using the left string . First, the characters in are mapped to string embeddings . Then, these embeddings are encoded to create a hidden representation , representing the context of the left-side string . In the following, the attention map is computed by applying a dot product between the hidden representation and the visual features , transformed by learned weight matrices and . The formulation is as follows:
@@
where represents a predefined token encoding the left-side string. The attention map is used to weight the visual features , producing a feature corresponding to character . After processing through the classifier , the predicted class probabilities for is obtained to calculate the cross-entropy loss, where is learnable weights.
@@
The weight of the attention map records the relevance of to visual features , and moreover, represents the context of string . So only when the visual model incorporates the context from into the visual features of the target character , the attention map can maximize the relevance between and visual features of that character, thus accurately highlighting the corresponding position of character , as shown in Fig. 3. A similar process can be applied to the right-side string , where the corresponding attention map and visual feature contribute to the prediction . By leveraging the above scheme during training, SGM effectively guides the visual model in integrating linguistic context into visual features. Consequently, even when SGM is not used during inference, the linguistic context can still be maintained and enhancing the accuracy of CTC models.
Note that although SGM is a decoder-based module, during inference it has been discarded and SVTRv2 becomes a purely CTC model. In contrast, previous methods, such as VisionLAN [47] and LPV [58], despite modeling linguistic context using visual features, still rely on attention-based decoders to activate linguistic information during inference, a process that is incompatible with CTC models.